# Unifying Multimodal Large Language Model Capabilities and Modalities via Model Merging

### v1

@@ Abstract
While foundation models update slowly due to resource-intensive training requirements, domain-specific models evolve between updates. Model merging aims to combine multiple expert models into a single, more capable model, thereby reducing storage and serving costs while supporting decentralized model development. Despite its potential, previous studies have primarily focused on merging visual classification models or Large Language Models (LLMs) for code and math tasks. Multimodal Large Language Models (MLLMs), which extend the capabilities of LLMs through large-scale multimodal training, have gained traction. However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. In this paper, (i) we introduce the model merging benchmark for MLLMs, which includes multiple tasks such as VQA, Geometry, Chart, OCR, and Grounding, providing both LoRA and full fine-tuning models. Moreover, we explore how model merging can combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. (ii) We implement 10 model merging algorithms on the benchmark. Furthermore, we propose a novel method that removes noise from task vectors and robustly optimizes the merged vector based on a loss defined over task vector interactions, achieving an average performance gain of 2.48%. (iii) We find that model merging offers a promising way for building improved MLLMs without requiring data training. Our results also demonstrate that the complementarity among multiple modalities outperforms individual modalities. All code and checkpoints are publicly available here.
@@ 1 Introduction
Foundation models experience slow development cycles due to resource-intensive training requirements, while domain-specific models continuously improve during interim periods [21]. Various developers release their fine-tuned models on open-source communities such as Hugging Face. Model merging [85] aims to combine multiple expert models into a unified model with multiple capabilities. This approach reduces storage and serving costs through model reuse, while supporting decentralized development by enabling independent contributors to build models that can later be merged. Despite its potential, previous studies [4, 35, 86] have primarily focused on merging visual classification models across multiple datasets to extract representations, or merging Large Language Models (LLMs) specifically for code and math tasks.
Multimodal Large Language Models (MLLMs), which extend LLMs with broader capabilities through large-scale multimodal training, have gained traction. Model merging offers a cost-effective way to
@@
Figure 1: Unifying the capabilities or modalities of MLLMs from open-source communities via model merging, which is cost-effective.
@@
combine fine-tuned MLLMs with task-specific skills into a unified model (see Fig. 1). However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. Specifically, AdaMMS [19] proposes an unsupervised hyper-parameter selection method, but can only merge two MLLMs at a time. For example, merging LLaVA-OneVision-Qwen [43] into Qwen2-VL [77] on the Qwen2 architecture. UQ-Merge [65] treats each fine-tuning dataset in LLaVA-v1.5 [48] as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline.
@@
In this paper, we introduce a model merging benchmark for MLLMs, which includes diverse tasks such as VQA, Geometry, Chart, OCR, and Grounding. For each task, we collect comprehensive public datasets with at least 100k samples to ensure effective supervised fine-tuning, and select corresponding benchmarks to evaluate distinct capabilities. We derive an upper bound on the error between the merged model and expert models, proving that merging performance is influenced by the learning rate and iterations, which control the extent of parameter drift. Smaller parameter changes lead to easier merging. We choose two types of vision-language models: InternVL2.5 and Qwen2-VL, providing both LoRA and full fine-tuning checkpoints. Moreover, most existing MLLMs specialize in dual modalities, and incorporating new modality encoders requires re-training on new modality-text data. Generating high-quality multimodal instruction data is resource-consuming [36]. Therefore, we explore how model merging can efficiently combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. This offers a data-free way to reuse and integrate modality-specific encoders into a unified LLM.
@@
Based on our benchmark, we conduct an in-depth comparison and analysis of state-of-the-art merging methods in capability and modality merging settings. Furthermore, we propose a novel merging method that improves the task vector (i.e., the parameter change between fine-tuned and base models) optimization. This method optimizes the merged model based on a loss defined over task vector interactions and applies low-rank approximations to reduce redundant noise, achieving the best results. Combining multiple MLLMs without additional data training, the merged model can even outperform expert MLLMs in their respective capabilities and mixture data training. We also find that merging methods effectively integrate inputs from multiple modalities, outperforming models trained on individual modalities, thus emphasizing the complementary nature of modal information.
@@
  • Benchmark: We introduce the model merging benchmark with detailed categorization of MLLM capabilities, and explore how model merging can effectively combine different modalities of MLLMs. We hope this benchmark will help the model merging community better evaluate the generalizability of their methods.
  • Methodology: We further propose a novel merging method that effectively removes noise from task vectors and improves the robustness of merged vector optimization. Ablation studies show that it achieves an average performance gain of 2.48%.
  • Experiments: We conduct comprehensive experiments and discussions on our benchmark. The empirical results demonstrate that model merging potentially surpasses mixture training, serving as a way for omni model alignment, while offering a scalable solution for developing MLLMs with reduced computational cost and time.
@@ 2 Related Work
Model merging has emerged as a cost-effective approach to developing improved models by combining multiple expert models to leverage their complementary capabilities [85, 3]. These expert models typically share a common base model, with specialization achieved through fine-tuning on distinct datasets. This approach offers a flexible and modular method for post-training MLLMs and facilitates the integration of new capabilities into top-performing models. Current research on model merging falls into two primary categories: static merging and dynamic merging. Static merging compresses multiple models into a single standard-sized model without adding additional computation or memory overhead. Dynamic merging (aka MoE-like methods) [74, 32, 52] requires the dynamic loading of task-specific modules based on test inputs, involving training routers or prior knowledge. The storage parameters for dynamic merging are larger.
Static merging can be further divided into data-free methods and test-time adaptation methods. Data-free methods merge fine-tuned models without requiring additional data. We categorize these methods into four groups: (i) Linear interpolation methods that perform arithmetic operations on task vectors [82, 35, 26, 11]; (ii) Sparsification-based methods that reduce redundancy in task vectors [84, 93, 31]; (iii) SVD-based methods that identify and exploit the low-rank features of task vectors [24, 56, 15, 70]; and (iv) Optimization-based methods that optimize task vectors via gradient descent [80, 14]. Test-time adaptation [88, 87, 17] assumes access to unlabeled test datasets, which can be considered a form of transductive learning.
@@
Recently, several works have attempted model merging for MLLMs, but with different objectives. VL-merging [71] merges modality-specific pre-trained models to create modality-agnostic models, evaluating their effectiveness through fine-tuning on downstream tasks (e.g., image-text retrieval and image classification). VisionFuse [12] employs task arithmetic to merge LLMs with concatenated visual encoder outputs, primarily focusing on enhancing MLLMs’ visual capabilities. DAMC [7] composes MLLMs across image, audio, video, and point cloud modalities while reducing modal interference through parameter decoupling.
Several approaches similar to ours aim to merge multiple MLLMs to improve multi-task performance. AdaMMS [19] proposes an unsupervised hyperparameter selection method for model merging. However, it requires generating responses for each candidate hyperparameter, making it time-consuming and assuming test set availability during merging. Furthermore, it can only merge two models at a time. For example, merging LLaVA-OneVision-Qwen into Qwen2-VL on the Qwen2 architecture, or merging LLaVA-v1.5 into CogVLM-Chat on the LLaMA architecture. UQ-Merge [65] considers uncertainty quantification on text and vision inputs to examine MLLM prediction confidence, requiring unlabeled test sets to calculate prediction and determine merging sequence. This approach is computationally expensive as it measures uncertainty across all candidate models and repeatedly evaluates merged models to find optimal combinations. UQ-Merge treats each fine-tuning dataset in LLaVA-v1.5 [48] as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline. In contrast, our benchmark collects more comprehensive data with clearer MLLM task divisions for fine-tuning, and we propose a data-free method that requires no hyperparameter search.
@@ 3 Rethinking Model Merging
Model merging aims to integrate multiple fine-tuned models, all derived from a base model $\boldsymbol{\theta}_{0}$, into a unified model that consolidates knowledge from diverse sources. Given $n$ fine-tuned models denoted as $\boldsymbol{\theta}_{1},\cdots,\boldsymbol{\theta}_{n}$, the objective is to produce a single merged model $\boldsymbol{\theta}_{m}$ that effectively inherits the capabilities of all individual models.
@@ 3.1 Merging Baselines
We categorize existing merging methods into four groups and briefly introduce each.
Linear interpolation methods: Weight Averaging [82] simply averages the weights of models fine-tuned on different tasks. Task Arithmetic [35] computes task vectors $\boldsymbol{\tau}_{i}=\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{0}$ for individual tasks and sums them to form a multi-task vector $\boldsymbol{\tau}_{m}=\sum_{i=1}^{n}\boldsymbol{\tau}_{i}$. This vector is then scaled by a coefficient $\lambda$ and added to the base model $\boldsymbol{\theta}_{0}$ to obtain the merged model.
Sparsification-based methods: Ties-Merging [84] combines steps like trimming, parameter sign determination, and disjoint merging to produce a merged task vector $\boldsymbol{\tau}_{m}$. The final model is defined as $\boldsymbol{\theta}_{m}=\boldsymbol{\theta_{0}}+\lambda\boldsymbol{\tau}_{m}$, where $\lambda$ is tuned using the validation set. DARE [93] randomly drops redundant task vectors and rescales the remaining ones to mitigate parameter interference.
SVD-based methods: TSV Merging [24] quantifies task-specific feature overlap in weight space by measuring the singular task interference of $\boldsymbol{\tau}_{i}$. It then reduces task interference through decorrelation, which is mathematically formulated as an orthogonal Procrustes problem. The method seeks orthogonal matrices $V_{\bot}$ and $U_{\bot}$ to reconstruct the parameters of the merged model. Iso-C [56] proposes an isotropic merging framework that flattens the singular value spectrum of task matrices, and enhances alignment between singular components of task-specific and merged matrices.
Optimization-based methods: WUDI Merging [14] proves that task vectors $\boldsymbol{\tau}$ form an approximate linear subspace of the fine-tuning data $\boldsymbol{x}$. This property allows the implicit utilization of training data information through task vectors alone. They define layer-wise interference between the merged vector and task vector as $\boldsymbol{\tau}_{m,l}-\boldsymbol{\tau}_{i,l}$ for task $i$ at layer $l$. To optimize the merged vector $\boldsymbol{\tau}_{m,l}$, they minimize this interference $(\boldsymbol{\tau}_{m,l}-\boldsymbol{\tau}_{i,l})\boldsymbol{x}_{i}$ with respect to task input $\boldsymbol{x}_{i}$. Leveraging the linear subspace relationship, they substitute the transpose of $\boldsymbol{\tau}_{i,l}$ for $\boldsymbol{x}_{i,l}$:
$\min_{\boldsymbol{\tau}_{m,l}}\mathcal{L}_{l}=\sum_{i=1}^{n}\frac{1}{\left\|\boldsymbol{\tau}_{i,l}\right\|_{F}^{2}}\left\|(\boldsymbol{\tau}_{m,l}-\boldsymbol{\tau}_{i,l})(\boldsymbol{\tau}_{i,l})^{\top}\right\|_{F}^{2}.$ (1)
@@ 3.2 Parameter Changes during Fine-tuning Matter
Model merging exhibits sensitivity to task vectors $\boldsymbol{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base model). Several studies [35, 93, 45] demonstrate that less intensive fine-tuning can yield superior merging performance, even when these models achieve lower accuracy on their respective tasks. In App. B, we conduct experiments on the impact of fine-tuning steps on merging performance, and observe that performance tends to rise initially and then decline. This counter-intuitive finding suggests that higher-performing expert models do not necessarily produce better merging outcomes. Fine-tuned models tend to converge around the base model in parameter space [61, 16]. When constructing our benchmark, we deliberately reduce parameter changes by lowering the learning rate while ensuring performance improvements on specific tasks. We analyze the upper bound of the loss incurred by model merging:
@@
Let $\boldsymbol{\theta}_{0}\in\mathbb{R}^{d}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\boldsymbol{\tau}_{i}=\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\boldsymbol{\tau}_{m}=\sum_{j}\alpha_{j}\boldsymbol{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then: $\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (2) This indicates that both the learning rate and iterations influence model merging results. Please refer to App. A for detailed assumptions and proofs.
@@
We examine the weight magnitude distribution of task vectors across our benchmark (see Fig. 2 (a-b)). Our analysis reveals that InternVL2.5, which undergoes full fine-tuning, exhibits a right-skewed distribution with parameter values concentrated around 1e-3. In contrast, Qwen2-VL, fine-tuned using LoRA, displays a multi-modal distribution with parameter values fluctuating between 1e-4 and 1e-5. Both models demonstrate distinct magnitude distribution patterns across different tasks. We also compute the normalized Frobenius norm of parameters (i.e., divided by the number of parameters). As shown in Fig. 2 (c-d), the Frobenius norm varies significantly across tasks and layers, which presents challenges that we will address in our approach. The small task vector magnitudes suggest that fine-tuned models and base models exist in adjacent regions of the loss landscape with linear connectivity [83], facilitating effective model merging [63, 75].

### v2

# OptMerge: Unifying Multimodal LLM Capabilities and Modalities via Model Merging

@@ Abstract
Foundation models update slowly due to resource-intensive training, whereas domain-specific models evolve rapidly between releases. Model merging seeks to combine multiple expert models into a single, more capable model, reducing storage and serving costs while supporting decentralized development. Despite its potential, previous studies have primarily focused on merging visual classification models or Large Language Models (LLMs) for code and math tasks. Recently, Multimodal LLMs (MLLMs) that extend LLMs through large-scale multimodal training have gained traction. However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. In this paper, (i) we introduce a model merging benchmark for MLLMs, which includes multiple tasks such as VQA, Geometry, Chart, OCR, and Grounding, studying both LoRA and full fine-tuning models. Moreover, we explore how model merging can combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. (ii) We implement 10 model merging algorithms on the benchmark. Furthermore, we propose a novel method that removes noise from task vectors and robustly optimizes the merged vector based on a loss defined over task vector interactions, achieving an average performance gain of 2.48%. (iii) We find that model merging offers a promising way for building improved MLLMs without requiring training data. Our results also demonstrate that the complementarity among multiple modalities outperforms individual modalities. All code and checkpoints are publicly available here.
@@ 1 Introduction
Figure 1: Unifying the capabilities or modalities of MLLMs from open-source communities via model merging, which is a data-free, cost-effective post-hoc method.
Foundation models experience slow development cycles due to resource-intensive training requirements, while domain-specific models continuously improve during interim periods (Fang et al., 2025). Various developers release their fine-tuned models on open-source communities such as Hugging Face. Model merging (Yadav et al., 2024) aims to combine multiple expert models into a unified model with multiple capabilities. This approach reduces storage and serving costs through model reuse, while supporting decentralized development by enabling independent contributors to build models that can later be merged. Despite its potential, previous studies (Akiba et al., 2025; Ilharco et al., 2023; Yang et al., 2024b) have primarily focused on merging visual classification models across multiple datasets to extract representations, or merging Large Language Models (LLMs) specifically for code and math tasks.
@@
Recently, Multimodal Large Language Models (MLLMs) that extend LLMs with broader capabilities through large-scale multimodal training have gained traction. Model merging offers a cost-effective way to combine fine-tuned MLLMs with task-specific skills into a unified model (see Fig.˜1). However, there lacks a benchmark for model merging research that clearly divides the tasks for MLLM training and evaluation. Specifically, AdaMMS (Du et al., 2025) proposes an unsupervised hyper-parameter selection method, but can only merge two MLLMs at a time. For example, merging LLaVA-OneVision-Qwen (Li et al., 2025a) into Qwen2-VL (Wang et al., 2024b) on the Qwen2 architecture. UQ-Merge (Qu et al., 2025) treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline.
@@
In this paper, we introduce a model merging benchmark for MLLMs, which includes diverse specialized models such as VQA, Geometry, Chart, OCR, and Grounding. For each corresponding task, we collect comprehensive public datasets with at least 100k samples to ensure effective supervised fine-tuning, and select corresponding benchmarks to evaluate distinct capabilities. We derive an upper bound on the error between the merged model and expert models, proving that merging performance is influenced by the learning rate and iterations, which control the extent of parameter drift. Smaller parameter changes sacrifice task performance but lead to easier merging. We choose two types of vision-language models: InternVL2.5 and Qwen2-VL, providing both LoRA and full fine-tuning checkpoints. Moreover, most existing MLLMs specialize in dual modalities, and incorporating new modality encoders requires re-training on new modality-text data. Generating high-quality multimodal instruction data is resource-consuming (Jiang et al., 2024). Therefore, we explore how model merging can efficiently combine different modalities (e.g., vision-language, audio-language, and video-language models), moving toward the Omni-language model. This offers a data-free way to reuse and integrate modality-specific encoders into a unified LLM.
@@
Based on our benchmark, we conduct an in-depth comparison and analysis of state-of-the-art merging methods in capability and modality merging settings. Furthermore, we propose a novel merging method that improves the task vector (i.e., the parameter change between fine-tuned and base models) optimization. OptMerge optimizes the merged model based on a loss defined over task vector interactions and applies low-rank approximations to reduce redundant noise, achieving the best results. Combining multiple MLLMs without requiring data, the merged model can even outperform expert MLLMs in their respective capabilities and mixture data training. We also find that merging methods effectively integrate inputs from multiple modalities, outperforming models trained on individual modalities, thus emphasizing the complementary nature of modal information.
@@
  • Benchmark: We introduce the first model merging benchmark that provides a fine-grained categorization of MLLM capabilities, and evaluates how merging integrates multiple modalities. We train expert models for each task and publicly release their weights and code. This benchmark is designed to help the model merging community better evaluate the generalizability of their methods.
  • Methodology: We further propose a simple yet effective method, OptMerge, which removes noise from task vectors and enhances the robustness of merged vector optimization. Ablation studies show an average performance improvement of 2.48%.
  • Experiments: We conduct comprehensive experiments and analyses on our benchmark. Our empirical results suggest that model merging can outperform mixture training, offering a viable path to omni-model alignment and a scalable approach to developing MLLMs with reduced computational cost and training time.
@@ 2 Related Work
Model merging has emerged as a cost-effective approach to developing improved models by combining multiple expert models to leverage their complementary capabilities (Yadav et al., 2024; Ahmadian et al., 2024). These expert models typically share a common base model, with specialization achieved through fine-tuning on distinct datasets. This approach offers a flexible and modular method for post-training MLLMs and facilitates the integration of new capabilities into top-performing models. Current research on model merging falls into two primary categories: static merging and dynamic merging. Static merging compresses multiple models into a single standard-sized model without adding additional computation or memory overhead. Dynamic merging (aka MoE-like methods) (Tang et al., 2024b; Huang et al., 2024; Lu et al., 2024b) requires the dynamic loading of task-specific modules based on test inputs, involving training routers or prior knowledge. The storage parameters for dynamic merging are larger.
Static merging can be further divided into data-free methods and test-time adaptation methods. Data-free methods merge fine-tuned models without requiring additional data. We categorize these methods into four groups: (i) Linear interpolation methods that perform arithmetic operations on task vectors (Wortsman et al., 2022; Ilharco et al., 2023; Goddard et al., 2024; Chen et al., 2025); (ii) Sparsification-based methods that reduce redundancy in task vectors (Yadav et al., 2023; Yu et al., 2024; He et al., 2025); (iii) SVD-based methods that identify and exploit the low-rank features of task vectors (Gargiulo et al., 2025; Marczak et al., 2025; Choi et al., 2024; Stoica et al., 2025); and (iv) Optimization-based methods that optimize task vectors via gradient descent (Wei et al., 2025b; Cheng et al., 2025). Test-time adaptation (Yang et al., 2024d; c; Daheim et al., 2024) assumes access to unlabeled test datasets, which can be considered a form of transductive learning.
@@
Recently, several works have attempted model merging for MLLMs, but with different objectives. VL-merging (Sung et al., 2023) merges modality-specific models to create modality-agnostic models, evaluating their effectiveness through fine-tuning on downstream tasks (e.g., image classification). VisionFuse (Chen et al., 2024d) employs task arithmetic to merge LLMs with concatenated visual encoder outputs, primarily focusing on enhancing MLLMs’ visual capabilities. Firstly, UnIVAL (Shukor et al., 2023) proposes a study on multimodal model merging via weight interpolation of models trained on different multimodal tasks, showing their benefits in particular for generalization. DAMC (Chen et al., 2024a) composes MLLMs across image, audio, video, and point cloud modalities while reducing modal interference through parameter decoupling.
Several approaches similar to ours aim to merge multiple MLLMs to improve multi-task performance. AdaMMS (Du et al., 2025) proposes an unsupervised hyperparameter selection method for model merging. However, it requires generating responses for each candidate hyperparameter, making it time-consuming and assuming test set availability during merging. Furthermore, it can only merge two models at a time. For example, merging LLaVA-OneVision-Qwen into Qwen2-VL on the Qwen2 architecture, or merging LLaVA-v1.5 into CogVLM-Chat on the LLaMA architecture. UQ-Merge (Qu et al., 2025) considers uncertainty quantification on text and vision inputs to examine MLLM prediction confidence, requiring unlabeled test sets to calculate prediction and determine merging sequence. This approach measures uncertainty across all candidate models and repeatedly evaluates merged models to find optimal combinations. UQ-Merge treats each fine-tuning dataset in LLaVA-v1.5 (Liu et al., 2024a) as a separate task without categorization of MLLM capabilities, fine-tuning the base model for each dataset and using LLaVA-v1.5 as the mixture training baseline. In contrast, our benchmark collects more comprehensive data with clearer MLLM task divisions for fine-tuning, and we propose a data-free method that requires no hyperparameter search.
@@ 3 Rethinking Model Merging
In Sec.˜3.1, we begin by introducing common model merging algorithms. In Sec.˜3.2, we revisit empirical findings from prior work, and provide a theoretical explanation of the relationship between model fine-tuning and merging performance. Building on this, we analyze the statistical properties of our benchmark, demonstrating both its validity and the challenges it presents.
@@ 3.1 Merging Baselines
Model merging aims to integrate multiple fine-tuned models, all derived from a base model $\bm{\theta}_{0}$, into a unified model that consolidates knowledge from diverse sources. Given $n$ fine-tuned models denoted as $\bm{\theta}_{1},\cdots,\bm{\theta}_{n}$, the objective is to produce a single merged model $\bm{\theta}_{m}$ that effectively inherits the capabilities of all individual models. We categorize merging methods into four groups.
Linear interpolation methods: Weight Averaging (Wortsman et al., 2022) simply averages the weights of models fine-tuned on different tasks. Task Arithmetic (Ilharco et al., 2023) computes task vectors $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$ for individual tasks and sums them to form a multi-task vector $\bm{\tau}_{m}=\sum_{i=1}^{n}\bm{\tau}_{i}$. This vector is scaled by a coefficient $\lambda$ and added to the base model $\bm{\theta}_{0}$ to obtain the merged model.
Sparsification-based methods: Ties-Merging (Yadav et al., 2023) combines steps like trimming, parameter sign determination, and disjoint merging to produce the $\bm{\tau}_{m}$. The final model is defined as $\bm{\theta}_{m}=\bm{\theta_{0}}+\lambda\bm{\tau}_{m}$, where $\lambda$ is tuned using the validation set. DARE (Yu et al., 2024) randomly drops redundant task vectors and rescales the remaining ones to mitigate parameter interference.
SVD-based methods: TSV Merging (Gargiulo et al., 2025) quantifies task-specific feature overlap in weight space by measuring the singular task interference of $\bm{\tau}_{i}$. It then reduces task interference through decorrelation. The method seeks orthogonal matrices $V_{\bot}$ and $U_{\bot}$ to reconstruct the parameters of the merged model. Iso-C (Marczak et al., 2025) proposes an isotropic merging framework that flattens the singular value spectrum of task matrices, and enhances alignment between singular components of task-specific and merged matrices.
Optimization-based methods: WUDI Merging Cheng et al. (2025) proves that task vectors $\bm{\tau}$ form an approximate linear subspace of the fine-tuning data $\bm{x}$. This property allows the implicit utilization of training data information through task vectors alone. They define layer-wise interference between the merged vector and task vector as $\bm{\tau}_{m,l}-\bm{\tau}_{i,l}$ for task $i$ at layer $l$. To optimize the merged vector $\bm{\tau}_{m,l}$, they minimize this interference $(\bm{\tau}_{m,l}-\bm{\tau}_{i,l})\bm{x}_{i,l}$ with respect to data $\bm{x}_{i,l}$. Leveraging the linear subspace relationship, they substitute the transpose of $\bm{\tau}_{i,l}$ for $\bm{x}_{i,l}$:
$\min_{\bm{\tau}_{m,l}}\mathcal{L}_{l}=\sum_{i=1}^{n}\frac{1}{\left\|\bm{\tau}_{i,l}\right\|_{F}^{2}}\left\|(\bm{\tau}_{m,l}-\bm{\tau}_{i,l})(\bm{\tau}_{i,l})^{\top}\right\|_{F}^{2}.$ (1)
This formulates model merging as a data-free optimization problem over parameters. Using the Adam optimizer, we obtain the merged vector $\bm{\tau}_{m,l}$, which minimizes interference with task vectors on multiple tasks, i.e., the hidden activation satisfies $(\bm{\theta}_{0,l}+\bm{\tau}_{m,l})\bm{x}_{i,l}\approx(\bm{\theta}_{0,l}+\bm{\tau}_{i,l})\bm{x}_{i,l}$.
@@ 3.2 Parameter Changes during Fine-tuning Matter
Model merging exhibits sensitivity to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base model). Several studies (Yu et al., 2024; Li et al., 2025b) demonstrate that less intensive fine-tuning can yield superior merging performance, even when these models achieve lower accuracy on their respective tasks. In App. B.1, we conduct experiments on the impact of fine-tuning steps on merging performance, and observe that performance tends to rise initially and then decline. This counter-intuitive finding suggests that higher-performing expert models do not necessarily produce better merging outcomes. Fine-tuned models tend to converge around the base model in parameter space (Merlin et al., 2023; Chung et al., 2024). When constructing our benchmark, we reduce parameter changes by lowering the learning rate while ensuring performance improvements on specific tasks. We analyze the upper bound of the loss incurred by model merging:
@@
Let $\bm{\theta}_{0}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (2)
This indicates that both the learning rate and iterations influence model merging results. Please refer to App. A for detailed assumptions and proofs.
Remark. Theorem˜3.1 provides the first theoretical explanation of how model fine-tuning affects merging performance. It supports previous empirical observations: fine-tuned models in current benchmarks typically remain within the same basin near the base model. For example, merging Qwen2.5-Math (Yang et al., 2024a) and Qwen2.5-Coder (Hui et al., 2024) yields poor performance. This is likely due to excessive post-training, which causes significant parameter drift. This insight suggests that poor merging results may not reflect algorithmic flaws, but rather issues with the fine-tuned models. When selecting models from Hugging Face, it’s helpful to choose fine-tuned models that are better suited for merging to minimize multi-task degradation.
@@
We examine the weight magnitude distribution of task vectors across our benchmark (see Fig.˜2 (a-b)). Our analysis reveals that InternVL2.5, which undergoes full fine-tuning, exhibits a right-skewed distribution. In contrast, Qwen2-VL, fine-tuned using LoRA, displays a multi-modal distribution. This is due to the low-rank nature and scaling factors of LoRA, which constrain the task vectors to be linear combinations in a reduced subspace, causing them to cluster along a few dominant magnitudes. Both models demonstrate distinct magnitude distribution patterns across different tasks. We also compute the normalized Frobenius norm of parameters (i.e., divided by the number of parameters). As shown in Fig.˜2 (c-d), the Frobenius norm varies significantly across tasks and layers, which presents challenges that we will address in our approach. The small task vector magnitudes suggest that fine-tuned models and base models exist in adjacent regions of the loss landscape with linear connectivity (Wu et al., 2023), facilitating effective model merging (Ortiz-Jimenez et al., 2023).